# Process-Supervised Reward Models for Verifying Clinical Note Generation: A Scalable Approach Guided by Domain Expertise

### v2

@@ 4.1 Main Results
Figure 3 presents a comparison of model performance across various evaluation tasks, including both ID and ODD scenarios. Across all metrics, our proposed PRM method consistently outperforms the vanilla ORM, as well as both Gemini Pro 1.0 and 1.5. Our model functions as either an ORM or PRM, differing in the aggregation strategy from step-wise to note-level scoring, as detailed in Section 3.4. Notably, our ORM implementation significantly outperforms the vanilla ORM, which predicts only the overall correctness of a generation without learning to estimate step-level reward signals. This suggests that even when functioning as an ORM and relying solely on the final end-of-note score, the ability to predict step-level correctness during training remains essential for achieving strong performance. In verification tasks, PRMs perform comparably to our ORMs when using the same model checkpoint, aligning with prior findings Beeching et al. ; Snell et al. (2024). However, in preference-based evaluations (A-Prefer), PRM surpasses our ORM, achieving 56.2% accuracy compared to 51.2%.
PRM excels in identifying gold-reference samples from error-containing ones in ID datasets, achieving near-perfect accuracy. Therefore, our primary analysis focuses on its generalization performance in OOD settings. Despite the inherent challenges of OOD tasks, PRM exhibits strong generalization, achieving 98.8% accuracy on the A-Verify task. Among all tasks, A-Prefer proves to be the most challenging, reflecting the complexity of physician preferences. Here, PRM achieves 56.2% accuracy in selecting physician-preferred notes among three candidates, significantly outperforming all baselines.
Finally, our results suggest that Gemini Pro models establish a strong baseline when used as LLM-as-a-judge with chain-of-thought reasoning. Notably, Gemini Pro 1.5 exhibits significant improvements over Gemini Pro 1.0 across all metrics, reflecting marked advancements in model capability.
Table 3: Ablation studies on loss functions and data selection. We evaluated different computational strategies for cross-entropy loss and assessed Gemini Pro-filtered data, both with and without paraphrases. The best performing model used notes-only loss and the full dataset with paraphrases. PRM and ORM results are reported from the checkpoint with the highest PRM A-Prefer performance. Numbers represent percentages of accuracy. Bold values indicate the highest PRM and ORM accuracy in each ablation study for ODD tasks. In ID tasks, multiple models achieved perfect accuracy.
@@ 4.2 Ablation Studies
We further investigated the impact of various loss functions, all derived from the same cross-entropy loss framework but differing in the selection of tokens included in the loss computation. The formal definitions of these loss functions are detailed in Appendix A. Restricting the loss computation to only the score label tokens resulted in a marked degradation of performance. Including losses over all special tokens (step tokens and score label tokens) produced performance similar to the vanilla approach but with slightly worse outcomes in verification tasks.
Notably, masking out losses from dialogue tokens while restricting the loss computation to the notes led to a positive impact on performance, particularly in the A-Prefer task. These findings suggest that it is important for the PRM to learn to generate notes (based on the dialogue) while simultaneously learning to predict reward labels in a step-by-step manner. This observation may be influenced by the fact that our PRM did not undergo task-specific supervised fine-tuning prior to reward model training. Masking dialogue losses likely helps prevent overfitting, as dialogues often include lengthy, repetitive content across samples from the same case.
@@ 5 Conclusions and Future Work
In this work, we proposed a methodology for developing PRMs for open-ended text generation in the clinical domain—guided by domain expertise yet fundamentally simple. We demonstrated the effectiveness of PRM in verifying clinical notes step-by-step and selecting physician-preferred notes via Best-of-N. Informed by key design principles rooted in real-world clinician expertise, our method is scalable and leverages synthetic data to automate process-supervised annotations for PRM training. Notably, our work is among the first to demonstrate the effectiveness of PRMs for open-ended generative tasks beyond mathematics and coding. These promising results suggest the potential for PRMs to extend beyond the clinical domain into a wide range of diverse generative tasks.
While our PRM achieves the highest performance in predicting physician preferences, it is important to note that no current method achieves consistently strong results. This highlights the complexity of the task and suggests that additional factors influencing alignment with physicians remain unexplored, representing valuable opportunities for future research. Lastly, the application of PRM for inference-time compute and scaling (e.g., MCTS) or step-by-step reinforcement learning is an exciting direction for future exploration.
@@ B.1 Training of PRM
All models were trained using either four 80GB NVIDIA A100 or H100 GPUs, employing DeepSpeed ZeRO Stage 3 optimization and the AdamW optimizer. We adopted Huggingface’s Transformers library and utilized its Trainer module. A limited search was conducted over learning rates [3e-4, 1.5e-4, 3e-5], with 3e-5 selected as the optimal value. The training process utilized consistent hyperparameters across all models, including a global batch size of 16, BF16 precision training, a warm-up phase of 50 steps, and a linear learning rate scheduler. All experiments were trained using one epoch of data.
@@ B.3 Use PRMs at Inference Time
At inference time, we first used a regular expression to transform the LLM-generated clinical notes into the JSON format defined in our steps. We replace score label tokens with a placeholder token when concatenating the note with the dialogue to form the input context. We then performe a standard forward pass with the input context and obtained the softmax probability of the special “$+$” token at each step position, which represents the PRM score for that step.
@@ B.5 Physician Reader Study
When selecting top-scoring samples for physician review, we first attempt to select those in which the probability of each step’s “$+$” score label exceeds that of its “--” score label. This approach mimics the strategy used in Lightman et al. (2023). If no such samples exist, we select the top-scoring sample from all candidates.
@@ D.2 Training PRMs with an Open-Source Framework
The core of our methodology lies in designing “errors” and structuring “steps” to align with real-world clinical documentation needs. While we opted to use the proprietary Gemini model to generate synthetic errors due to its strong performance and ease of API access, our methodology remains model-agnostic. These synthetic errors could just as effectively be generated using other powerful open-source models, such as LLaMA-3.1 405B Dubey et al. (2024) or Deepseek-R1 Guo et al. (2025) (given adequate GPU resources), provided that the critical decisions regarding error types are made beforehand. Furthermore, our model training is already based on the open-source LLaMA model. Lastly, while we used Gemini Pro to generate synthetic cases for PRM development, PRMs outperform both Gemini Pro 1.0 and 1.5, indicating that the proprietary model does not impose an upper bound on PRM performance within our methodology.
D.3 Focusing on the Assessment and Plan of Clinical Note
Building on recent research on ambient clinical charting and the real-world feedback Wang et al. (2024a), we concentrate solely on the Assessment and Plan (A&P) section of the clinical note generated from patient-doctor dialogues for the following reasons: (1) The A&P section is arguably the most critical and time-intensive part of a clinical note, as it encapsulates key clinical reasoning and medical decision-making. (2) After domain-specific adaptation, the “Subjective” section produced by LLMs is generally of high quality and nearly indistinguishable from human-written content, whereas notable gaps persist in the A&P section. (3) Other sections, such as the “Objective” or “Physical Exam,” are deemed less significant, as they are often imported from the electronic medical record or align less effectively with the physician workflow.
@@ Appendix E Additional Related Work
In this exciting new paradigm of large-scale RL, we believe that PRMs remain effective for the following reasons: (1) As a simple yet powerful approach, PRMs serve as an effective means for Best-of-N filtering, guiding test-time computation. (2) In general domains where a verifiable answer is difficult to define, PRMs provide a reliable single reward signal when functioning as ORMs. As demonstrated in our study, this approach outperforms the standard ORM implementation. Moreover, recent research further highlights the potential of PRMs in guiding test-time scaling, offering both effectiveness and computational efficiency Liu et al. (2025); Cheng et al. (2025).
@@ Appendix G PRM Examples
We presented additional example notes along with their PRM scores in Appendix Table 10 and 11.
@@ 8. JSON Schema:
Appendix Table 8: Prompts to Gemini Pro for preferred note selection.

### v3

@@ 4.1 Main Results
Comparison with Baselines
We present the main results in Figure 3. Our PRM consistently outperforms all baselines across all tasks, including the vanilla ORM, Gemini Pro 1.5, and SOTA reasoning models. Our analysis focuses on PRM performance in OOD settings (A-Verify and A-Prefer), given its near-perfect accuracy in the two ID verification tasks (Dialogue-G and A-Validate). Despite the inherent challenges of OOD tasks, PRM exhibits strong generalization, achieving 98.8% accuracy on the A-Verify task (compared to 93.8% for Gemini Pro 1.5). Among all tasks, A-Prefer proves to be the most challenging, reflecting the complexity of physician preferences. Here, PRM achieves 56.2% accuracy in selecting physician-preferred notes (compared to 50.0% for Gemini Pro 1.5). We discuss the explanation for the strong OOD performance in Section C.3. Importantly, while we utilized Gemini Pro 1.5 to generate synthetic data for PRM training, our results demonstrate that PRMs significantly outperform Gemini Pro 1.5 itself. This suggests that the proprietary model used for data generation doesn’t impose an inherent upper bound on PRM performance within our methodology.
PRM vs. ORM
Our model can function as either an ORM or a PRM, differing in the aggregation strategy from step-wise to note-level scoring, as detailed in Section 3.4. In verification tasks, PRMs perform comparably to our ORMs when using the same model checkpoint, aligning with prior findings Beeching et al. ; Snell et al. (2024). However, in preference-based evaluations (A-Prefer), PRM surpasses our ORM, achieving 56.2% accuracy compared to 51.2%.
Comparison with Vanilla ORM
Interestingly, our ORM implementation significantly outperforms the vanilla ORM, which predicts only the overall correctness of a generation without learning to estimate step-level reward signals. This suggests that even when functioning as an ORM and relying solely on the end-of-note score, learning to predict step-level correctness during training is essential for achieving strong performance.
Performance of Proprietary Models
As expected, the most advanced reasoning models are better suited for the challenging A-Prefer task, with o3-mini delivering the strongest baseline performance. Notably, newer-generation models within each family show clear performance improvements (e.g., Gemini Pro 1.5 outperforms Gemini Pro 1.0, and o3-mini surpasses o1).
@@ 4.2 Ablation Studies
We further investigated the impact of various loss functions, all derived from the same cross-entropy loss framework but differing in the selection of tokens included in the loss computation. The formal definitions of these loss functions are detailed in Appendix A. Restricting the loss computation to only the score label tokens resulted in a marked degradation of performance. Including losses over all special tokens (i.e., step tokens and score label tokens) produced performance similar to the vanilla approach but with slightly worse outcomes in verification tasks.
Notably, masking out losses from dialogue tokens while restricting the loss computation to the notes led to a positive impact on performance, particularly in the A-Prefer task. These findings suggest that it is important for the PRM to learn to generate notes (based on the dialogue) while simultaneously learning to predict reward labels in a step-by-step manner. This observation may be influenced by the fact that our PRM did not undergo task-specific supervised finetuning prior to reward model training. Masking dialogue losses likely helps prevent overfitting, as dialogues often include lengthy, repetitive content across samples from the same case.
@@ 5 Conclusions and Future Work
In this work, we introduce a methodology for developing PRMs for open-ended text generation in the clinical domain. Our PRM achieved SOTA performance in verifying clinical notes step-by-step and selecting physician-preferred notes via Best-of-N. Our work is among the first to demonstrate the effectiveness of PRMs for open-ended generative tasks beyond mathematics and coding. Notably, our framework is generalizable; it integrates LLMs and domain knowledge to produce process-supervised data at scale, and can be easily applied to other generative tasks across various domains.
While our PRM achieves the highest performance in predicting physician preferences, it is important to note that no current method achieves strong results. This suggests that additional factors influencing alignment with physicians remain unexplored, representing valuable opportunities for future research. Lastly, the application of PRMs for inference-time compute and scaling (e.g., MCTS) or step-by-step reinforcement learning is an exciting direction for future exploration.
6 Acknowledgments
This research was supported by NSF awards SCH-2205289. The funder played no role in the study design, data collection, analysis, and interpretation of data, or the writing of this manuscript.
7 Ethical Considerations
Our work leveraged an LLM to generate synthetic data for training PRMs to verify clinical notes. In considering real-world deployment, several considerations are critical for minimizing harm in this clinical context. The scope of errors in synthetic data should be expanded beyond the four categories examined here to enable more comprehensive evaluation of clinical notes across diverse contexts. In addition, large-scale validation of PRM outputs is essential to identify unintended consequences such as overlooked errors or reward hacking. Finally, safeguards must be implemented to prevent synthetic data from introducing or amplifying biases, for instance when certain error types are disproportionately associated with specific disease populations, which could inadvertently reinforce inequities in documentation or care delivery.
@@ B.1 Training of PRM
All models were trained using either four 80GB NVIDIA A100 or H100 GPUs, employing DeepSpeed ZeRO Stage 3 optimization Rasley et al. (2020) and the AdamW optimizer Loshchilov and Hutter (2017). We adopted Huggingface’s Transformers library and utilized its Trainer module Wolf et al. (2019). A limited search was conducted over learning rates [3e-4, 1.5e-4, 3e-5], with 3e-5 selected as the optimal value. The training process utilized consistent hyperparameters across all models, including a global batch size of 16, BF16 precision training, a warm-up phase of 50 steps, and a linear learning rate scheduler. All experiments were trained using one epoch of data.
@@ B.3 Use PRMs at Inference Time
At inference time, we first use a regular expression to transform the LLM-generated clinical notes into the JSON format defined by our steps. We replace score label tokens with a placeholder token when concatenating the note with the dialogue to form the input context. We then perform a standard forward pass with the input context and obtain the softmax probability of the special “++” token at each step position, which represents the PRM score for that step.
@@ B.5 Physician Reader Study
When selecting top-scoring samples for physician review, we first attempt to select those in which the probability of each step’s “++” score label exceeds that of its “-” score label. This approach mimics the strategy used in Lightman et al. (2023). If no such samples exist, we select the top-scoring sample from all candidates.
@@ D.2 Training PRMs with an Open-Source Framework
The core of our methodology lies in designing “errors” and structuring “steps” to align with real-world clinical documentation needs. While we opted to use the proprietary Gemini model to generate synthetic errors due to its strong performance and ease of API access, our methodology is model-agnostic. These synthetic errors could just as effectively be generated using other powerful open-source models, such as LLaMA-3.1 405B Dubey et al. (2024) or Deepseek-R1 Guo et al. (2025) (given adequate GPU resources), provided that the critical decisions regarding error types are made beforehand. Furthermore, our model training is based on the open-source LLaMA model. Lastly, while we used Gemini Pro to generate synthetic cases for PRM development, PRMs outperform both Gemini Pro 1.0 and 1.5, indicating that the proprietary model does not impose an upper bound on PRM performance within our methodology.
D.3 Assessment and Plan of Clinical Note
Building on recent research on ambient clinical charting and the real-world feedback Wang et al. (2024a), we concentrate solely on the “Assessment and Plan” (A&P) section of the clinical note in our work for the following reasons: (1) The A&P section is arguably the most critical and time-intensive part of a clinical note, as it encapsulates key clinical reasoning and medical decision-making. (2) After domain-specific adaptation, the “Subjective” section produced by LLMs is generally of high quality and nearly indistinguishable from human-written content, whereas notable gaps persist in the A&P section. (3) Other sections are considered less significant, as they are often imported directly from the electronic medical record (e.g., “Objective”) or fit less naturally into the physician’s workflow during ambient scribing (e.g., “Physical Exam”).
D.4 Considerations in defining "steps" for clinical note
When defining the "steps" for clinical note, we took into account the diverse real-world requirements of clinical documentation from various stakeholders. For example, the problem list within a note plays a critical role in billing processes for insurance companies and also serves as a key reference for daily clinical communication, warranting special attention Chowdhry et al. (2017). In addition, we drew on existing guidelines and best practices for clinical note writing, which explicitly emphasize the value of structured, problem-oriented documentation Li et al. (2018); DeParle (2000).
@@ Appendix E Additional Related Work
In this exciting new paradigm of large-scale RL, we believe that PRMs remain effective for the following reasons: (1) As a simple yet powerful approach, PRMs serve as an effective means for Best-of-N filtering, guiding test-time computation. (2) In general domains where a verifiable answer is difficult to define, PRMs provide a reliable single reward signal when functioning as ORMs. As demonstrated in our study, this approach outperforms the standard ORM implementation. Moreover, recent research highlights the potential of PRMs in guiding test-time scaling, offering both effectiveness and computational efficiency Liu et al. (2025); Cheng et al. (2025).
@@ Appendix G PRM Examples
We presented additional example notes along with their PRM scores in Appendix Table 10 and $1$.
@@ 8. JSON Schema:
Appendix Table 8: Prompts to Gemini Pro 1.5 for preferred note selection.